SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (03/25/08)

Attendees: Su Dong, Rahmat Rahmat, Debbie Bard, Kim Hojeong.

Meeting Information

 Date      :  March 25th, 2008
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

  • Paged 3/23 1pm: L1 DCT-GLT & EMT-GLT transmission errors more than 0.3%. => power-cycle crates solved the problem.
    These errors are related to the rate of L1 trigger bursts with many events came 
    very close to each other in time. If there is some indication that L1 rate is 
    fluctuating more than usual, especially if there are more spikes in time history, 
    then this is not surprising. I can see the current background e.g. from the lost 
    particle at phi~200 degrees is worse than usual. 

EMT

DCT

GLT

Online

L3

  • New L3 Lumi applications installed by Rainer (3/23 9am):
    • [09:43] [Rainer] Installed a new set of the Level 3 luminosity applications, including the DHP end-of-run utilities that write luminosity and other run information into the logbook, as well as the Level 3 PCA server, which serves all Level 3 data to EPICS. The luminosity number will now come out corrected according to the offline calibration discussed recently in the luminosity HyperNews. As of last Friday, the numbers have already been corrected by an update script, retroactively from the beginning of Run 7; the new installed software automates that change, and any future changes in the correction. For the EPICS part of the change to become active, the L3TPcaServer will have to be restarted using the run control proxy button. I asked the shift crew to do that at their convenience. (It can be done independent of data taking.)
    • [10:13] [Rainer] I see that the L3TPcaServer was restarted, however it reran the version from the release. I assume for the change to take effect the proxy itself will have to be restarted also. There should be an option for that in the menu that pops up with the L3TPcaServer proxy button. Thanks.

DQM

  • Rahmat saw false FCT offlines run 80915-16 (March 23rd 7:45pm) had about 100,000 number of errors. => These runs are cosmics. The FCT/GLT line false on/off logic was hardwired to 15 unprescaled (and unmasked) lines for colliding beam config. For cosmic runs with only two lines masked on, it will essentially complain on every event.

Trickle injection monitoring

Trigger Configuration

  • During above-Y(4S) scan, we will use current Y(2S) configurations.

AOB

  • Su Dong will be gone for 2 weeks. Starting next Tuesday.

This page was created by Kim Hojeong; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: March 25th, 2008.